# Rethinking Score Distillation as a Bridge Between Image Distributions

### v1

@@ 2.1 Background
Diffusion models define a forward “noising" process that degrades data samples $\mathbf{x}$ gradually from the image distribution to noised samples $\mathbf{z}_{t}$, and eventually the i.i.d. Gaussian distribution [23, 60]. This process is indexed by timesteps $t$, where $t=1$ indexes the full Gaussian noise distribution and $t=0$ indexes the data distribution. A diffusion model, parameterized by $\phi$, is then trained to reverse this encoding process, iteratively transforming the noise distribution into the data distribution with the score-matching objective:
@@
where $y$ is a conditioning text prompt, and $\alpha_{t}$ and $\sigma_{t}$ are hyperparameters from the predefined noise schedule.
@@
where $f(\mathbf{x},t)$ and $g(t)$ are pre-defined schedule coefficients. This PF-ODE can be solved deterministically [61], mapping a noise sample to its corresponding data sample through the reverse process and the opposite through the forward process (inversion). This cycle-consistent conversion between image and latent representations is important in establishing dual diffusion implicit bridges.
@@ 2.2 Optimization with Diffusion Model Approximates a Dual Schrödinger Bridge
Many generative vision tasks involve optimizing corrupted images to the image manifold. For example, in 3D generation, a 3D representation like NeRF is optimized to render natural images matching a prescribed text prompt. Methods like SDS enable this by using a pre-trained diffusion model as a prior. We propose formulating such optimization problems as solutions to an instantiation of a Schrödinger Bridges Problem (SBP). SBP finds cost-optimal paths between a source image distribution $p_{\text{src}}$ and a target image distribution $p_{\text{tgt}}$ [66, 14]. Optimizing a parametrized image toward the natural image distribution can be cast as finding the optimal paths between the current optimized image(s) and the natural image distribution. Instead of solving this problem directly, which would require training a generative model from scratch  [38, 14, 10], we show that pre-trained diffusion models can be exploited as an optimizer that approximates the path. Further, the gradient computed by the existing score distillation methods can be viewed as the first-order approximation of this path. This formulation is illustrated in Figure 1
@@
$\mathbf{x}_{\theta,t}=\alpha_{t}\mathbf{x}+\sigma_{t}\mathbf{\epsilon}$
@@
    First-order approximation error. Instead of performing full PF-ODE simulations, the single-step noising and prediction are less accurate and can induce errors. Recent work ISM [34] can be interpreted as reducing this error with a multi-step simulation to obtain $\mathbf{x}_{\theta,t}$.
@@ 2.3 Analyzing Existing Score Distillation Methods
Experimentally, we produce very similar results at high CFG with or without the non-dominant terms. We argue that SDS should be interpreted through the dominant term, which fits within our analysis. Under this interpretation, the unconditional distribution approximates the distribution of $\textbf{x}_{\theta}$ poorly, instead representing images of any identity with low contrast and geometric artifacts. Figure 1(b) illustrates the effect of a poor approximation. The bridge from the unconditional to conditional distribution leads to the characteristic oversaturation and smoothing of SDS results.
@@ 2.4 Mitigating Source Distribution Mismatch with Textual Descriptions
To describe the artifacts produced by SDS, we append the descriptors “, oversaturated, smooth, pixelated, cartoon, foggy, hazy, blurry, bad structure, noisy, malformed” and drop the descriptors of the high-quality generation. Note that in all of our generation experiments, the description of $y_{\text{src}}$ is fixed as above. We explored searching for other prompts but did not find that variations in these descriptions made a big difference.
@@ 3 Experiments
In this section, we test our proposed method on several generation problems where SDS is adopted. We compare against SDS and other task-specific baselines. Note that our goal is not to show another state-of-the-art text-to-3D generation method, but to verify our findings, where the proposed score distillation approach based on textual description efficiently improves the results by mitigating the source distribution mismatch error. We first perform a thorough experiment in a controlled setting on zero-shot text-to-image generation. Then, we compare it on text-guided NeRF optimization to SDS and VSD and evaluate the painting-to-real image translation task against image editing baselines. Please see more qualitative results, as well our method’s application to optical illusion generation and 3D-sketch-to-real task, in the appendix.
@@ 3.1 Zero-Shot Text-to-Image Generation with Score Distillation
To verify our analysis of existing SDS variants and the proposed method, we perform text-to-image generation by optimizing an image of size $64\times 64\times 4$ in the Stable Diffusion latent space [68, 28]. The benefit of choosing image generation as the evaluation task is that its generation quality has the least confounding variables among other tasks. (e.g., in text-to-3D, many designs like regularizations [75], initialization [35], 3D representations [9, 65, 72, 64], and 2D prior models [55, 40, 39, 47, 76] could affect the final quality.)
@@ 3.2 Text-guided NeRF Optimization
We now evaluate the text-to-3D generation problem, where we intentionally aim to exclude variables that could affect the generation quality other than the score distillation methods. We use the ThreeStudio [19] repository to optimize a NeRF with settings tuned for ProlificDreamer stage 1 (NeRF optimization) [68]. Note that we do not perform stages 2 and 3, i.e. geometry fine-tuning and texture refinement. Specifically, we initialize the NeRF with the method proposed by Magic3D [35], use the regularization losses on the sparsity and opacity, and optimize for 25K steps. We adopt the native SDS and VSD guidance implementations for comparison.
@@
We also perform a quantitative evaluation and user study on the NeRFs optimized based on 31 different text prompts. Note that this number is similar to the choice of existing works on the text-to-3D task [34, 32, 15]. However, different from these works that ignore the confounding 3D variables that contribute to the generation quality, we disentangle this by isolating the score distillation method as the only comparison variable. We follow these works to evaluate the generation quality with CLIP [48]. We report the CLIP similarity in Table 2. Our method consistently outperforms SDS and achieves comparable results with VSD. In addition, in a user study consisting of 37 users, shown pairwise comparisons of rotating 3D renders (i.e., comparisons of our result and a random choice of VSD or SDS, with the prompt: “For a text-to-3D system, given the prompt [p], which result would you be happiest with?”), our results were chosen in 75.7% of all responses. We also show more results in the Appendix.
Figure 5: Painting-to-Real comparison. We compare our gradient in optimization to image restoration and image-conditional generation baselines. While SDEdit produces convincing textures, its difficult to find a strength value that balances structure and quality. Other baselines fail to reproduce natural image quality, while our method produces the best combination of quality and faithfulness.
3.3 Painting-to-Real.
@@
4 Discussion
@@
As we have shown that reducing the distribution mismatching error can significantly improve the generation quality of the score distillation optimization, it is natural to ask whether one can also reduce the first approximation error, induced by linear bridge estimation, to improve the results further. Several recent studies including SDI [41] and ISM [34] can be viewed as mitigating this error by replacing the single-step estimation with multi-step estimation. Instead of performing multiple PF-ODE steps, one can solve the entire PF-ODE path to recover the dual bridge and estimate the endpoint of the bridge $\psi_{0,\text{tgt}}$ that is coupled with $\psi_{0,\text{src}}$. In this way, we obtain the most accurate gradient direction with little approximation error $\mathbf{\epsilon}_{\text{SBP}}^{*}=w\cdot(\psi_{0,\text{tgt}}-\psi_{0,\text{src}})$.
@@
However, solving the inversion ODE is not trivial [27]. We noticed that the inversion can exaggerate the distribution mismatch error and cause the optimization to get stuck at a local optimal at the beginning of the optimization. Instead, the high variance of the single-step methods often shows more robustness to the input image. Therefore, we first perform the single-step score distillation optimization to obtain reasonable results before moving to solving the full bridge. We find that in text-to-2D, such a method can produce high-quality results closer to the DDIM sampling results, as demonstrated by a COCO-FID score of $55.65$, which is better than VSD results. However, the same trend does not fully transfer to the text-to-3D experiments. We observe that it typically introduces additional artifacts and makes the optimization less stable. We leave the best way of leveraging this gradient as a future research exploration.
5 Related Work
5.1 Score Distillation Sampling
Although modalities like 3D, 4D, sketch, and vector graphics (SVGs) lack the large-scale, diverse, and high-quality datasets needed to train a domain-specific diffusion model, previous works find it useful to exploit image or video as a proxy modality [26, 16]. By computing the gradient on a proxy representation with a pretrained model, optimization in the target modality is viable with differentiable mappings, e.g. differentiable rasterization [33] for SVGs or differentiable rendering [44] for 3D objects and scenes. The seminal method, Score Distillation Sampling (SDS) [46], first proposed to apply a pretrained text-to-image diffusion model for text-to-3D generation. However, it requires a high classifier-free guidance weight and, therefore, suffers from artifacts such as over-saturation and over-smoothing. Recent works have built upon SDS to adapt it for editing tasks [30, 20, 45, 29] or more broadly improve over the original SDS formulation [28, 1, 68, 75, 74, 76]. NFSD [28] and LMC-SDS [1] inspect the individual components of the SDS gradient and propose methods to rectify the high guidance weights. However, the over-saturation problem is mitigated but not fully resolved. VSD [68] formulates the problem as particle-based variational inference and proposes to train a LoRA [24] on the fly to estimate the score of proxy distribution. We presented a new framework that allows rethinking all the variants under the same lens. This framework also motivates a method that improves the quality of SDS without losing efficiency.
5.2 Visual Content Generation with SDS
Since SDS was developed for text-to-3D generation, it has also been adopted to generate various other visual content such as SVGs [18, 71], sketches [70], texture [43, 6, 7, 8, 73], typography [25], dynamic 4D scenes [2, 58, 37] and illusions [5]. Among these applications, text-to-3D has been the most active research direction. In addition to designing better distillation sampling methods [68, 75, 28], prior work has also studied the underlying 3D neural representations [72, 64, 35, 9] and leveraging multiview data to improve the 3D consistency [55, 40, 39, 47, 76]. We note that these explorations are orthogonal to our study and should be able to work jointly with our method. In this paper, we looked into existing applications like text-based NeRF optimization, painting-to-real, and illusion generation. We also propose a new AR application called 3D sketch-to-real.
6 Closing Remarks
@@
We present an analysis that formulates the use of a pre-trained diffusion model in an optimization framework as seeking an optimal transport between two distributions. Under this lens, we analyze SDS variants with a unified framework. We also develop a simple approach based on textual descriptions that work comparably well to the best-performing approach, VSD, without its significant computational burden. However, neither approaches have yet to achieve the quality and diversity of images generated by the reverse process. We hope that our analysis enables the development of a more sophisticated solution that can one day achieve the same quality and diversity as the reverse process in an optimization framework. Combining our proposed method with multi-step approximations like ISM [34] or schedules like DreamFlow [32] could mitigate the first-order approximation error and further improve the efficiency, which is an interesting future research direction. With the rise of high-quality video diffusion models, we anticipate that the question of how to effectively use such models as a prior in various problems will become even more important.
Potential Social Impacts
@@ Acknowledgment.
We thank Matthew Tancik, Jiaming Song, Riley Peterlinz, Ayaan Haque, Ethan Weber, Konpat Preechakul, Amit Kohli and Ben Poole for their helpful feedback and discussion. This project is supported in part by a Google research scholar award and IARPA DOI/IBC No. 140D0423C0035. The views and conclusions contained herein are those of the authors and do not represent the official policies or endorsements of these institutions.
@@ Appendix A Additional Experimental Setup
For CSD, we follow the original paper [74] to use $w_{1}=w_{2}=40$ at the initialization steps and anneal $w_{2}=0$ within the first $500$ steps. We use $s=100$ for SDS and $s=7.5$ for NFSD and VSD, which are consistent with the best practice. We use $s=40$ and $w=25$ for our method. And we optimize with $\mathbf{\epsilon}_{\text{SDS}}$ loss for $500$ iterations and then switch to $\mathbf{\epsilon}_{\text{ours}}$ for the rest of $2,000$ iterations. For all the methods, we use a learning rate of $0.01$, and we use a learning rate of $41e-$ 4 to train the LoRA in VSD.

### v2

@@ 1 Introduction
2 Related Work
Score Distillation Sampling
Modalities like 3D, 4D, sketch, and vector graphics (SVGs) lack the large-scale, diverse, and high-quality datasets needed to train a domain-specific diffusion model. In these domains, previous works explore exploiting image or video as a proxy modality [26, 16]. By computing the gradient on a proxy representation with a pretrained model, optimization in the target modality is viable with differentiable mappings, e.g. differentiable rasterization [33] for SVGs or differentiable rendering [44] for 3D objects and scenes. The seminal method, Score Distillation Sampling (SDS) [48], first proposed to apply a pretrained text-to-image diffusion model for text-to-3D generation. However, it requires a high classifier-free guidance weight and, therefore, suffers from artifacts such as over-saturation and over-smoothing. Recent works have built upon SDS to adapt it for editing tasks [30, 20, 46, 29] or more broadly improve over the original SDS formulation [28, 1, 70, 77, 76, 78]. NFSD [28] and LMC-SDS [1] inspect the individual components of the SDS gradient and propose methods to rectify the high guidance weights. However, the over-saturation problem is mitigated but not fully resolved. VSD [70] formulates the problem as particle-based variational inference and proposes to train a LoRA [24] on the fly to estimate the score of proxy distribution. We present a new framework that allows rethinking all the variants under the same lens. This framework also motivates a method that improves the quality of SDS without losing efficiency.
Visual Content Generation with SDS
Since SDS was developed for text-to-3D generation, it has also been adopted to generate various other visual content such as SVGs [18, 73], sketches [72], texture [43, 6, 7, 8, 75], typography [25], 3D bodies [45], dynamic 4D scenes [2, 60, 37] and illusions [5]. Among these applications, text-to-3D has been the most active research direction. In addition to designing better distillation sampling methods [70, 77, 28], prior work has also studied the underlying 3D neural representations [74, 66, 35, 9] and leveraging multiview data to improve the 3D consistency [57, 40, 39, 49, 78]. We note that these explorations are orthogonal to our study and should be able to work jointly with our method. In this paper, we look into existing applications like text-based NeRF optimization, painting-to-real, and illusion generation. We also propose a new AR application called 3D sketch-to-real.
@@ 3.1 Background
Diffusion models define a forward “noising" process that degrades data samples $\mathbf{x}$ gradually from the image distribution to noised samples $\mathbf{z}_{t}$, and eventually the i.i.d. Gaussian distribution [23, 62]. This process is indexed by timesteps $t$, where $t=1$ indexes the full Gaussian noise distribution and $t=0$ indexes the data distribution. A diffusion model, parameterized by $\phi$, is then trained to reverse this encoding process, iteratively transforming the noise distribution into the data distribution with the following denoising objective:
@@
where $w(t)$ is a loss weighting function, $y$ is a conditioning text prompt, and $\alpha_{t}$ and $\sigma_{t}$ are hyperparameters from the predefined noise schedule.
@@
where $f(\mathbf{x},t)$ and $g(t)$ are pre-defined schedule parameters. This PF-ODE can be solved deterministically [63], mapping a noise sample to its corresponding data sample through the reverse process and the opposite through the forward process (inversion). This cycle-consistent conversion between image and latent representations is important in establishing dual diffusion implicit bridges.
@@ 3.2 Optimization with Diffusion Model Approximates a Dual Schrödinger Bridge
Many generative vision tasks involve optimizing corrupted images to the image manifold. For example, in 3D generation, a 3D representation like NeRF is optimized to render natural images matching a prescribed text prompt. Methods like SDS enable this by using a pre-trained diffusion model as a prior. We propose formulating such optimization problems as solutions to an instantiation of a Schrödinger Bridges Problem (SBP). SBP finds cost-optimal paths between a source image distribution $p_{\text{src}}$ and a target image distribution $p_{\text{tgt}}$ [68, 14]. Optimizing a parametrized image toward the natural image distribution can be cast as finding the optimal paths between the current optimized image(s) and the natural image distribution. Instead of solving this problem directly, which would require training a generative model from scratch  [38, 14, 10], we show that pre-trained diffusion models can be exploited as an optimizer that approximates the path. Further, the gradient computed by the existing score distillation methods can be viewed as the first-order approximation of this path. This formulation is illustrated in Figure 1.
@@
$\mathbf{x}_{\theta,t}=\alpha_{t}{\color[rgb]{0,0,0}\mathbf{x}_{\theta}}+\sigma_{t}\mathbf{\epsilon}$ (4)
@@
    First-order approximation error. Instead of performing full PF-ODE simulations, the single-step noising and prediction are less accurate and induce errors. Recent work ISM [34] can be interpreted as reducing this error with a multi-step simulation to obtain $\mathbf{x}_{\theta,t}$.
@@ 3.3 Analyzing Existing Score Distillation Methods
Experimentally, we produce very similar results at high CFG with or without the non-dominant terms. We argue that SDS should be interpreted through the dominant term, which fits within our analysis. Under this interpretation, the unconditional direction $\phi\left(\mathbf{x}_{\theta,t};\varnothing,t\right)$ approximates the source distribution of $\textbf{x}_{\theta}$ poorly, instead representing images of any identity with low contrast and geometric artifacts. Figure 1(b) illustrates the effect of a poor approximation. The bridge from the unconditional to conditional distribution leads to the characteristic oversaturation and smoothing of SDS results.
@@ 3.4 Mitigating Source Distribution Mismatch with Textual Descriptions
To describe the artifacts produced by SDS, we append the descriptors “, oversaturated, smooth, pixelated, cartoon, foggy, hazy, blurry, bad structure, noisy, malformed” and drop the descriptors of the high-quality generation. This description $y_{\text{src}}$ does not require hand-crafting based on problem domains—it is fixed across all shown examples and use cases. As shown in Appendix Figure A5, we explored searching for other prompts but did not find that variations in these descriptions made a big difference.
@@ 4 Experiments
In this section, we test our proposed method on several generation problems where SDS is adopted. We compare against SDS and other task-specific baselines. Note that our goal is not to show another state-of-the-art text-to-3D generation method, but to verify our findings, where the proposed score distillation approach based on textual description efficiently improves the results by mitigating the source distribution mismatch error. We first perform a thorough experiment in a controlled setting on text-to-image generation. Then, we compare it on text-guided NeRF optimization to SDS and VSD and evaluate the painting-to-real image translation task against image editing baselines. Please see more results in the appendix, including additional qualitative results and comparison, ablation studies and our method’s application to optical illusion generation and 3D-sketch-to-real task.
@@ 4.1 Zero-Shot Text-to-Image Generation with Score Distillation
To verify our analysis of existing SDS variants and the proposed method, we perform text-to-image generation by optimizing an image of size $64\times 64\times 4$ in the Stable Diffusion latent space [70, 28] (We explore other base models like MVDream [57] and SDXL [47] in Appendix Figure A3). The benefit of choosing image generation as the evaluation task is that its generation quality has the least confounding variables among other tasks. (e.g., in text-to-3D, many designs like regularizations [77], initialization [35], 3D representations [9, 67, 74, 66], and 2D prior models [57, 40, 39, 49, 78] could affect the final quality.)
@@ 4.2 Text-guided NeRF Optimization
We now evaluate the text-to-3D generation problem, where we intentionally aim to exclude variables that could affect the generation quality other than the score distillation methods. We use the ThreeStudio [19] repository to optimize a NeRF with settings tuned for ProlificDreamer stage 1 (NeRF optimization) [70]. Note that we do not perform stages 2 and 3, i.e. geometry fine-tuning and texture refinement. Specifically, we initialize the NeRF with the method proposed by Magic3D [35], use the regularization losses on the sparsity and opacity, and optimize for 25K steps. We adopt the native SDS and VSD guidance implementations for comparison. In Appendix Figure A4, we evaluate our methods with additional text-to-3D systems, including Fantasia3D [9], Magic3D [35] and CSD [76].
@@
We also perform a quantitative evaluation and user study on the NeRFs optimized based on 31 different text prompts. Note that this number is similar to the choice of existing works on the text-to-3D task [34, 32, 15]. However, different from these works that ignore the confounding 3D variables that contribute to the generation quality, we disentangle this by isolating the score distillation method as the only comparison variable. We follow these works to evaluate the generation quality with CLIP [50]. We report the CLIP similarity in Table 2. Our method consistently outperforms SDS and achieves comparable results with VSD. In addition, in a user study consisting of 37 users, shown pairwise comparisons of rotating 3D renders (i.e., comparisons of our result and a random choice of VSD or SDS, with the prompt: “For a text-to-3D system, given the prompt [p], which result would you be happiest with?”), our results were chosen in 75.7% of all responses.
Figure 5: Painting-to-Real comparison. We compare our gradient in optimization to image restoration and image-conditional generation baselines. While SDEdit produces convincing textures, it is difficult to find a strength value that balances structure and quality. Other baselines fail to reproduce natural image quality, while our method produces the best combination of quality and faithfulness.
4.3 Painting-to-Real
@@
5 Discussion on Solving the Linear Approximation Error
Table 3: Reducing first-order approximation error improves generation quality. Using full PF-ODE simulation (“Full-path”) to replace single-step prediction improves visual quality in all settings.
Approximate the bridge Single-step Full-path
Estimate source distribution Uncond. (SDS) Bridge LoRA (VSD) Uncond. Bridge LoRA
Zero-Shot COCO FID ($\downarrow$) 86.02 67.89 59.22 63.31 60.07 55.65
@@ 5 Discussion on Solving the Linear Approximation Error
As we have shown that reducing the distribution mismatching error can significantly improve the generation quality of the score distillation optimization, it is natural to ask whether one can also reduce the first-order approximation error, induced by linear bridge estimation, to improve the results further. Several recent studies, including SDI [41] and ISM [34], can be viewed as mitigating this error by replacing the single-step estimation with a multi-step estimation to an intermediate timestep. Under our framework, one can estimate the entire dual bridge by solving both PF-ODE paths. Specifically, via inversion, one can solve the PF-ODE path from $\psi_{0,\text{src}}$ to $x_{\theta,T}$, and then walk to the $\psi_{0,\text{tgt}}$ via sampling. In this way, it is possible to obtain the most accurate gradient direction with little approximation error $\mathbf{\epsilon}_{\text{SBP}}^{*}=w\cdot(\psi_{0,\text{tgt}}-\psi_{0,\text{src}})$. We refer to this approach as “full path”. Note that this resolves the linear approximation error, and it is independent of handling the source approximation error, which could be addressed via the discussed text description or LoRA.
@@
However, solving the inversion ODE is not trivial [27]. We noticed that the inversion can exaggerate the distribution mismatch error and cause the optimization to get stuck at a local optimum at the beginning of the optimization. Instead, the stochasticity of the single-step methods often shows more robustness to the input image. Therefore, we first perform the single-step score distillation optimization to obtain reasonable results and then switch to solving the full bridge. We also anneal the timestep endpoint of the bridge throughout the optimization. With this approach, we can now explore addressing both the first and second sources of error. The first source (linear approximation) has “full-path,” and the second source (source distribution mismatch error) has “Bridge” or “LoRA”. We find that using the “full-path” multi-step (mitigating linear approximation error) always outperforms the single-step methods, achieving a lower FID, as shown in Table 3. However, the same trend does not fully transfer to the text-to-3D experiments. We observe that solving the entire bridge typically introduces additional artifacts and makes the optimization less stable. We leave the best way of leveraging this gradient for future research exploration.
6 Conclusion
@@
We present an analysis that formulates the use of a pre-trained diffusion model in an optimization framework as seeking an optimal transport between two distributions. Under this lens, we analyze SDS variants with a unified framework. We also develop a simple approach based on textual descriptions that work comparably well to the best-performing approach, VSD, without its significant computational burden. However, neither approach has yet to achieve the quality and diversity of images generated by the reverse process. We hope that our analysis enables the development of a more sophisticated solution that can one day achieve the same quality and diversity as the reverse process in an optimization framework. Combining our proposed method with multi-step approximations like ISM [34] or schedules like DreamFlow [32] could mitigate the first-order approximation error and further improve the efficiency, which is an interesting future research direction. With the rise of high-quality video diffusion models, we anticipate that the question of how to effectively use such models as a prior in various problems will become even more important.
@@ Acknowledgment.
We thank Matthew Tancik, Jiaming Song, Riley Peterlinz, Ayaan Haque, Ethan Weber, Konpat Preechakul, Amit Kohli and Ben Poole for their helpful feedback and discussion. This project is supported in part by a Google research scholar award, IARPA DOI/IBC No. 140D0423C0035, and NSF grant No. IIS-2213335. The views and conclusions contained herein are those of the authors and do not represent the official policies or endorsements of these institutions.
@@ Appendix A Additional Experimental Setup
We use the stable-diffusion-v2-1-base model by default for our experiments if not specified. For CSD, we follow the original paper [76] to use $w_{1}=w_{2}=40$ at the initialization steps and anneal $w_{2}=0$ within the first $500$ steps. We use $s=100$ for SDS and $s=7.5$ for NFSD and VSD, which are consistent with the best practice. We use $s=40$ and $w=25$ for our method. And we optimize with $\mathbf{\epsilon}_{\text{SDS}}$ loss for $500$ iterations and then switch to $\mathbf{\epsilon}_{\text{ours}}$ for the rest of $2,000$ iterations. For all the methods, we use a learning rate of $0.01$, and we use a learning rate of $41e-$ 4 to train the LoRA in VSD.